### v1

# Hire Me or Not? Examining Language Model’s Behavior with Occupation Attributes

Abstract
With the impressive performance in various downstream tasks, large language models (LLMs) have been widely integrated into production pipelines, like recruitment and recommendation systems. A known issue of models trained on natural language data is the presence of human biases, which can impact the fairness of the system. This paper investigates LLMs’ behavior with respect to gender stereotypes, in the context of occupation decision making. Our framework is designed to investigate and quantify the presence of gender stereotypes in LLMs’ behavior via multi-round question answering. Inspired by prior works, we construct a dataset by leveraging a standard occupation classification knowledge base released by authoritative agencies. We tested three LLMs (RoBERTa-large, GPT-3.5-turbo, and Llama2-70b-chat) and found that all models exhibit gender stereotypes analogous to human biases, but with different preferences. The distinct preferences of GPT-3.5-turbo and Llama2-70b-chat may imply the current alignment methods are insufficient for debiasing and could introduce new biases contradicting the traditional gender stereotypes.
1 Introduction
Large language models (LLMs) have become well-known to public users due to their impressive performance across multiple tasks (Tan et al., 2023; Wang et al., 2023; Lee et al., 2023) that are scalable with model size (Kaplan et al., 2020). Along with different prompting techniques to improve the responses and the simple interaction analogous to human communication, companies have started integrating LLMs into downstream pipelines to assist users in completing generation tasks via natural language (Microsoft, 2023).
However, a known issue of language models (LMs) is the human biases traced back to the large training corpus (Bender et al., 2021; Blodgett et al., 2020; Nozza et al., 2022; Smith et al., 2022; Solaiman et al., 2019; Talat et al., 2022), which can impact the fairness of the downstream tasks (Rudinger et al., 2018; Zhao et al., 2018a; Stanovsky et al., 2019; Dev et al., 2020; Liang et al., 2021; He et al., 2021). Various methods have been proposed to mitigate human biases, for example, data augmentation using counterfactuals (Zhao et al., 2019; Maudslay et al., 2019; Zmigrod et al., 2019), adjusting model parameters (Lauscher et al., 2021; Garimella et al., 2021; Kaneko & Bollegala, 2021; Guo et al., 2022), and modify decoding step to decrease the harmful generations (Schick et al., 2021). Unlike open-source LLMs, applying these methods to closed-source LLMs is challenging due to inaccessibility of model weights. Additionally, even if one can access the model weights, fine-tuning LLMs to mitigate a certain human bias may introduce new biases (Van Der Wal et al., 2022), as demonstrated by prior works in embedding debiasing methods (Bordia & Bowman, 2019; Gonen & Goldberg, 2019; Nissim et al., 2020). Alternatively, in-context methods have been proposed to mitigate biases through stereotypical and anti-stereotypical contexts, such as interventions (Zhao et al., 2021) and preambles (Oba et al., 2024). Recent work has found that LLMs still exhibit gender biases even after removing explicit signals, such as co-occurrences of “female” and “nurse”, suggesting the measured bias is not necessarily relevant to explicit gender-associated words (Belém et al., 2023).
Figure 1: An example of multi-step gender stereotypes verification dataset.
In this work, we focus on gender stereotypes related to occupation. Particularly, we investigate LLMs’ behavior with the appearance of implicit neutral occupation-relevant attribtues. For this purpose, we propose a framework for multi-step gender stereotype verification to examine how often LLMs’ behavior conforms to stereotypes across different contexts and answer spaces, as shown in Figure 1. As human biases change along with time and environment (Kozlowski et al., 2020), we leverage the latest standard occupation classification taxonomy released by O*NET (Gregory et al., 2019) as the source of implicit neutral occupation-relevant attributes.
Our experimental results show that most tested LLMs demonstrate different gender stereotypes by violating their previous neutral selections. Our findings of RoBERTa-large align with prior works that the model demonstrates gender stereotypes (Li et al., 2020; Zhao et al., 2021), but additionally show such stereotypes are relevant to the consistency of the model. The results of GPT-3.5-turbo and Llama2-70b-chat show some gender stereotypes are analogous to humans and some contradict traditional stereotypes. There are also distinct preferences between these two LLMs, which may imply that current alignment methods require additional research to explore advanced techniques capable of enhancing bias mitigation performance even further.
2 Related Work
Repetitive co-occurrences between genders and certain occupations could perpetuate and be transmitted through natural language then forming gender biases, for example, male doctors and female nurses. Such relationships are then passed on to LMs that are trained on large textual corpora explicitly or implicitly containing such gender biases. Extensive literature has shown that gender biases exist in the input representations to pre-trained language models (PLMs) (Bolukbasi et al., 2016; Caliskan et al., 2017; Zhao et al., 2017; Garg et al., 2018; Zhao et al., 2018b; May et al., 2019; Swinger et al., 2019; Zhao et al., 2019; Chaloner & Maldonado, 2019; Bordia & Bowman, 2019; Tan & Celis, 2019; Zhao et al., 2020), and downstream tasks, for example, coreference resolution (Rudinger et al., 2018; Zhao et al., 2018a; Kurita et al., 2019), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019; Cho et al., 2019), textual entailment (Sap et al., 2020; Dev et al., 2020), and so on (Tatman, 2017; Kiritchenko & Mohammad, 2018; Park et al., 2018; Sheng et al., 2019; Lu et al., 2020).
Some recent works have focused on probing models’ behavior via alternating the input (Wallace et al., 2019; Gardner et al., 2020; Sheng et al., 2020; Emelin et al., 2021; Ye & Ren, 2021; Schick & Schütze, 2021; Oba et al., 2024), as well as via underspecified questions (Li et al., 2020; Zhao et al., 2021).
A range of recent works investigates human biases in LLMs. Acerbi & Stubbersfield (2023) use transmission chain-like methodology to reveal that ChatGPT-3 shows biases analogous to humans for stereotypical content over other content. Gupta et al. (2024) find that LLMs are deeply biased and suggest they manifest implicit stereotypical and often erroneous presumptions when taking on a persona. Wan et al. (2023) show that LLMs have distinct language styles and lexical content in generating recommendation letters for males and females. Belém et al. (2023) demonstrate that measured gender bias is not necessarily due to the explicit signals, suggesting the implicit factors that contribute to the biased behavior of LLMs. Kotek et al. (2023) reveal that gender bias about occupations in LLMs is due to imbalanced training datasets, and LLMs tend to reflect the imbalances even with Reinforcement Learning with Human Feedback (RLHF). Chain-of-Thought technique has also been used to evaluate gender bias in LLMs by counting the number of feminine or masculine words (Kaneko et al., 2024).
Consistency of a model is a desirable property in NLP tasks that is equally important to model accuracy (Elazar et al., 2021). There are many prior works exploring consistency of PLMs for question answering (Rajpurkar et al., 2016; Ribeiro et al., 2019; Alberti et al., 2019; Asai & Hajishirzi, 2020; Kassner et al., 2021), robust evaluation(Li et al., 2019), natural language inference (Camburu et al., 2018; 2020), and more (Du et al., 2019).
3 Multi-step Gender Stereotypes Verification
In this paper, we introduce Multi-step Gender Stereotype Verification that involves three consecutive steps providing different contexts of occupation-relevant attributes, stereotypical occupation titles, underspecified questions, and different answer spaces, as shown in Figure 2. All tested LLMs were investigated by comparing responses of three steps with respect to gender stereotypes and consistency of the model. Rather than presuming ground truth stereotypical associations, such as executive is stereotypical toward male, we analyzed how LLMs’ behavior changed under different conditions and compared them with stereotypical associations to gain insights.
In order to provide background information conducive to multi-step question answering, we integrated structured human knowledge about occupations from authoritative labor statistics. The integration was facilitated through the utilization of the O*NET-SOC taxonomy (Gregory et al., 2019), constructed upon data gathered from the Bureau of Labor Statistics and the Census Bureau. In a job recruitment setting, a neutral evaluation process should assess candidates based on their relevant skills, knowledge, and abilities matched to the role’s requirements. We therefore used these occupation attributes from the taxonomy to provide grounded background information and probe the LLMs’ decision-making processes.
3.1 Dataset Construction
Our dataset focuses on two subject categories: male and female, aiming to investigate potential gender biases regarding various occupation titles. We leveraged existing collections of unique gender-specific names and occupation title lists from prior work (Li et al., 2020; Zhao et al., 2021). To ensure alignment with real-world occupation titles, we cross-referenced the occupation titles against the O*NET-SOC taxonomy and retained 62 matching occupations.
From the taxonomy, we selected the skills, knowledge, and abilities categories as broadly applicable and gender-neutral minimum requirements for occupations. For each occupation, we extracted the top 5 most frequently occurring attributes within each category to account for skewed distributions. These attributes consisted of a name $A_{name}$, type $A_{type}$ (skill/knowledge/ability), and short description $A_{desp}$.
Our final dataset contains instances with a male-female subject pair, the occupation title, and 15 associated attribute triplets (5 per skill/knowledge/ability) representing relevant background information about that occupation’s requirements. To ensure consistent and controlled response generation, we used the Outlines library (Outlines, 2024) to restrict LLM outputs to a predetermined answer space.
Figure 2: Multi-rounds of questions
3.2 Extracting Background Attribute Associations
To begin, we probed the relevant attributes of each occupation (skills, knowledge, abilities) associated with each subject in a gender pair using LLMs. Following the format of the UnQover dataset (Li et al., 2020), each instance contained a base context paragraph $p$ and an underspecified question $q$:
Paragraph: [$s_{1}$] and [$s_{2}$] are neighbors.
Question: Who was a [$occupation$]?
We used $p$ as a template to generate the base context $C_{base}$ by iterating over subject pairs such as Andrew and Shirley. The underspecified question $q$ was then adjusted with specific attribute descriptions from the O*NET taxonomy:
Cbase: [$s_{1}$] and [$s_{2}$] are neighbors.
Question: [$a_{desp}$]. Does [$s_{1}$/$s_{2}$] have [$a_{name}$] [$a_{type}$]?
An example of the attribute question is listed below:
Given [$s_{1}$]=Andrew, [$s_{2}$]=Shirley, [$a_{name}$, $a_{type}$, $a_{desp}$]=[Mathematics, Knowledge, Mathematics is defined as …]
Cbase: Andrew and Shirley are neighbors.
Attribute Question: Mathematics is defined as … Does Shirley have Mathematics knowledge?
In Step 1, as illustrated in Figure 2, each LLM outputs Yes/No to indicate whether the subject possesses a particular attribute. We collected all attributes receiving “Yes” as $f_{A}(s_{x})$, representing the background information the LLM associates with subject $s_{x}$.
3.3 Assessing Individual Qualifications
Next, for Step 2 in Figure 2, we evaluated whether the LLM deems each subject individually qualified for the occupation based solely on their inferred background information $f_{A}(s_{x})$:
Q1: [$C_{base}$]. [$f_{A}(s_{x})$]. Is [$s_{x}$] qualified for [$occupation$]?
Each LLM outputted a binary Yes/No response $R_{binary}(s_{x})$, indicating its assessment of the subject’s qualifications given their associated attributes.
3.4 Comparing Subject Selections
Finally, in Step 3 as illustrated in Figure 2, we probed which subject the LLM favors when considering background information $f_{A}(s_{male})$ and $f_{A}(s_{female})$ for both subjects. We used two meaning-preserved variants:
Q2: [$C_{base}$]. [$f_{A}(s_{female})$]. [$f_{A}(s_{male})$]. Who is more qualified to be a/an [$occupation$]?
Q3: [$C_{base}$]. [$f_{A}(s_{female})$]. [$f_{A}(s_{male})$]. Who will be a/an [$occupation$]?
$Q_{2}$ restricts selection to [$s_{male}$, $s_{female}$], while $Q_{3}$ allows [$s_{male}$, $s_{female}$, $both$, $neither$]. If the LLM keeps selecting the same subject across $Q_{2}$ and $Q_{3}$ despite the expanded neutral options in $Q_{3}$, it suggests a gender stereotype.
The LLM outputs are denoted as $R_{single}(s_{x})$ for $Q_{2}$ and $R_{multi}(s_{x})$ for $Q_{3}$. The right part of Figure 2 shows the process.
3.5 Metric
A key aspect of our multi-step verification framework is the ability to systematically analyze both potential gender stereotypes and consistency in LLMs’ behaviors. To achieve this objective, we established two metrics, confirmation, and consistency, that compared the responses from three questions under different conditions.
3.5.1 Confirmation
Question pairs $Q_{2}$ and $Q_{1}$ investigate the LLM on the same subject, but differ in implicitly neutral contexts and answer spaces. $Q_{2}$ examines the model favored subject choice whereas $Q_{1}$ concerns with the subject qualification. Jointly, we are able to evaluate if the LLM shows biased behavior by selecting subject $s_{x}$ in $Q_{2}$ and violating the individual qualification assessment in $Q_{1}$ across the evaluation set:
$Confirmation(L,Q_{1},Q_{2},D_{eval})=\frac{1}{|D_{eval}|}\sum_{(s_{f},s_{m})\in D_{eval}}\Phi(L(Q_{2})==s_{x},L(Q_{1},s_{x})==Yes)$ (1)
where $L$ represents the LM and $\Phi(*)$ returns 1 if both conditions are met (LLM selected $s_{x}$ in $Q_{2}$ and also answered Yes that $s_{x}$ is qualified in $Q_{1}$), and 0 otherwise.
3.5.2 Consistency
The meaning-preserving question pairs $Q_{2}$ and $Q_{3}$ investigate the LLM on the same decision, but alter the answer space from [$s_{male}$, $s_{female}$] to [$s_{male}$, $s_{female}$, $both$, $neither$]. This enables us to evaluate if the LLM exhibits biased behavior by persistently favoring the same subject across $Q_{2}$ and $Q_{3}$ despite the additional neutral options in $Q_{3}$’s answer space:
$Consistency(L,Q_{1},Q_{2},D_{eval})=\frac{1}{|D_{eval}|}\sum_{(s_{f},s_{m})\in D_{eval}}\Phi(L(Q_{2}),L(Q_{3}))$ (2)
where $\Phi(*)$ outputs 1 if the responses to $Q_{2}$ and $Q_{3}$ are identical for the [$s_{male}$, $s_{female}$] subject pair, and 0 otherwise.
Taken together, a high score on Confirmation and a low score on Consistency would suggest an LLM exhibiting low gender bias in its occupational decision-making process.
4 Results
We evaluated the following three LLMs. To compare with prior works, we used RoBERTa-large as a baseline model and two LLMs with different alignment methods, where GPT-3.5-turbo uses Reinforcement Learning with Human Feedback (RLHF) and Llama2-70b-chat uses RLHF and Supervised Fine-Tuning (SFT). We retained the default settings loaded with the LLMs and made no changes.
  • RoBERTa-large (Liu et al., 2019) fine-tuned on SQuAD v2.0 (Rajpurkar et al., 2018)
  • GPT-3.5-turbo (OpenAI, 2021)
  • Llama2-70b-chat (Touvron et al., 2023)
Figure 3: Measurement between $Q_{1}$ and $Q_{2}$
Figure 3 displays the Confirmation metric, measured as whether the LLM’s $Q_{2}$ subject selection matches its $Q_{1}$ individual qualification assessment for that subject. Compared to RoBERTa-large, GPT-3.5-turbo exhibits higher variance in Confirmation, indicating greater fluctuation in its decision as additional subject background information is provided. Llama2-70b-chat demonstrates lower variance but with some outliers, indicating generally stable but occasional deviations from its own qualification judgments.
Figure 4: Measurement between $Q_{2}$ and $Q_{3}$
Figure 4 shows the Consistency metric which evaluates whether LLMs maintain consistent outputs across the meaning-preserving question pairs with different answer choices. Notably, both GPT-3.5-turbo and Llama2-70b-chat exhibit lower overall scores than RoBERTa-large. The Consistency score of GPT-3.5-turbo distribution is concentrated toward 0, indicating the model tends to modify its behavior when providing more neutral options. The score of Llama2-70b-chat is between RoBERTa-large and GPT-3.5-turbo.
Figure 5: Scatterplot of $Q_{1}Q_{2}$ vs. $Q_{2}Q_{3}$
As shown in Figure 5, analyzing Confirmation and Consistency jointly shows interesting patterns across LLMs. RoBERTa-large demonstrates a relatively linear relationship, where occupations with high Confirmation scores also have high Consistency scores. Its gender stereotypes appear to be more systematic, as additional background information does not significantly alter its behavior.
In contrast, GPT-3.5-turbo exhibits a nearly vertical Confirmation-Consistency pattern heavily concentrated at 0 Consistency. This suggests providing additional neutral information successfully mitigates gender stereotypes in many cases, but inconsistently compared to its initial qualification decisions.
The scatter of Llama2-70b-chat is focused on the top-left quadrant, with reasonably high Confirmation but low Consistency scores across occupations.
5 Discussion and Future Work
Finally, we would like to discuss these observations and where they could lead to understanding of LLMs.
The distributions of three LLMs’ joint confirmation and consistency scores in Figure 5 are quite intriguing. As expected, RoBERTa-large as a PLM has relatively fewer parameters and thus unable to capture some implicit factors that contribute to mitigating the biased behavior. GPT-3.5-turbo utilized RLHF to align its responses with human preference. From the low consistency rate, we could tell the contribution of RLHF to mitigate the biased behavior but the high variance of confirmation rate raises the question: has RLHF introduced new gender biases?
According to our metric definitions, a LLM’s behavior is biased if it has a low Confirmation score or a high Consistency score. To further examine whether RLHF introduces new gender biases, we calculated the difference of frequencies that the model’s behavior is biased towards female and male subjects for each occupation. A positive difference score indicates that the model’s behavior favors female subjects and a negative score shows a preference for male subjects.
Occupation Diff
politician 0.840
senator 0.815
piano player 0.590
violin player 0.590
film director 0.525
guitar player 0.520
doctor 0.320
poet 0.285
lawyer -0.235
plumber -0.260
janitor -0.290
butcher -0.330
driver -0.360
hunter -0.396
athlete -0.400
mechanic -0.620
pilot -0.660
Table 1: Occupation with high difference via GPT-3.5-turbo. Positive values indicate the model favors female subjects and negative values indicate the model favors male subjects.
Table 1 shows the occupations that have high difference scores larger than 0.2 from the results of GPT-3.5-turbo. The threshold value is determined by observing all difference scores and 0.2 is an explicit boundary. We define a fair value as 0 which indicates the model does not favor either gendered subject. Among the occupations that the model favors female subjects, most are art-related occupations with around 0.5 values. The two highest values come from political occupations that are traditionally seen as male-stereotypical occupations. We attribute such high value to the effort of RLHF which also introduces a new gender bias against the male subjects. Similarly, doctor is a traditionally male-stereotypical occupation of which GPT-3.5-turbo now favors the female subjects. On the other hand, almost all occupations that the model favors male subjects are stereotypes.
Occupation Diff
piano player 0.085
journal editor 0.075
film director 0.065
carpenter 0.050
manager 0.045
scientist 0.040
detective 0.035
writer 0.035
architect 0.030
assistant professor 0.030
hunter -0.045
violin player -0.045
bodyguard -0.050
model -0.090
pilot -0.090
athlete -0.130
Table 2: Occupation with high difference via Llama2-70b-chat. Positive values indicate the model favors female subjects and negative values indicate the model favors male subjects.
Table 2 shows the occupations that have high difference scores larger than 0.02 from the results of Llama2-70b-chat. As we already stated, the threshold value is determined by observing all difference scores and 0.02 is an explicit boundary. Occupations that the model favors female subjects are mixed with art-related occupations as well as science-related occupations, and the values are all close to 0 which can be ignored as stereotypes. Similarly, occupations that the model favors male subjects are from various domains and the values are negligible.
Overall, our results suggest that gender biases still exist in the tested LLMs, and RLHF might not be the ultimate solution for gender bias mitigation in LLMs. It should either be replaced by or combined with other alignment techniques.
5.1 Limitation
Our work is limited to investigating gender biases and stereotypes in English, a morphologically limited language. Recent studies have found gender biases existing in LLMs for different languages (Malik et al., 2022; Névéol et al., 2022; Kaneko et al., 2022; Anantaprayoon et al., 2023; Levy et al., 2023). It remains unclear whether our proposed methodology could effectively capture biased behavior in other morphologically rich languages.
Moreover, we focused solely on gender biases in this work. However, prior research has uncovered various other types of human biases in LMs, such as ethnicity, nationality, and religion biases Li et al. (2020); Zhao et al. (2021). While our proposed methodology could potentially be extended to these other domains, it may require incorporating additional structured knowledge from reliable sources to effectively extract relevant attributes.
Furthermore, we only considered gender-specific names in our work. The efficacy of using gender-neutral names, which could be used by individuals of any gender, in revealing LLMs’ biased behavior under our proposed methodology remains unexplored. Additionally, our work only addresses binary gender biases, whereas non-binary gender biases have also been explored in recent literature (Cao & Daumé III, 2020; Dev et al., 2021).
6 Conclusion
In this work, we proposed a multi-step gender stereotypes verification framework to investigate LLMs’ potentially biased behavior across different implicitly neutral contexts and answer spaces. Our methodology does not require access to LLMs’ weights, making it broadly applicable. Our carefully crafted dataset leverages a reliable taxonomy to provide up-to-date structured knowledge of occupation-relevant attributes. Additionally, we introduced two novel metrics, Confirmation and Consistency, to systematically evaluate both potential gender stereotypes and consistency in LLMs’ behavior.
Our experimental results show that LLMs still possess gender stereotypes analogous to human biases. Our findings for RoBERTa-large align with prior works. Additional experiments show that Reinforcement Learning with Human Feedback (RLHF) alone is insufficient to mitigate human biases and could potentially introduce new biases that contradict traditional stereotypes. The differences between the distributions of GPT-3.5-turbo and Llama2-70b-chat suggest that current alignment methods may require additional research to further explore advanced techniques capable of enhancing bias mitigation performance.
We urge caution in using LLMs in bias-sensitive domains without thorough testing to understand the potential impact and corresponding solutions for safe and equal treatment of all subjects. Our work provides a systematic framework for investigating and quantifying gender stereotypes in LLMs, contributing to future research in human bias mitigation and responsible AI development.